# Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

### v1

@@ Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes. However, humans usually perceive 3D objects or scenes at a higher level as a composition of parts or structures rather than points or voxels. Representing 3D as semantic parts can benefit further understanding and applications. We aim to solve part-aware 3D reconstruction, which parses objects or scenes into semantic parts. In this paper, we introduce a hybrid representation of superquadrics and 2D Gaussians, trying to dig 3D structural clues from multi-view image inputs. Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time. We incorporate parametric superquadrics in mesh forms into 2D Gaussians by attaching Gaussian centers to faces in meshes. During the training, superquadrics parameters are iteratively optimized, and Gaussians are deformed accordingly, resulting in an efficient hybrid representation. On the one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes. On the other hand, 2D Gaussians are incorporated to model the complex texture and geometry details, ensuring high-quality rendering and geometry reconstruction. The reconstruction is fully unsupervised. We conduct extensive experiments on data from DTU and ShapeNet datasets, in which the method decomposes scenes into reasonable parts, outperforming existing state-of-the-art approaches.
Index Terms:
Part-aware reconstruction, Hybrid representation, 2D Gaussians, Superquadrics.
{strip}[Uncaptioned image]
@@ I Introduction
3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which is different from human perceptions. Humans understand 3D scenes or objects as different semantic parts [3]. For example, when we see a scene, we are building high-level structural information such as scene graphs in our minds, not point clouds or detailed geometry. Thus, we propose to learn part-aware scene reconstruction to decompose the scenes into different individual semantic parts or shapes, which would be more helpful for tasks including scene manipulation/editing, scene graph generation, etc.
@@
There are prior methods [4, 5, 6, 7, 8, 9, 1, 10, 11, 12, 13] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios. While neural radiance field (NeRF) [14] has demonstrated the potential of reconstructing textured 3D scenes from multi-view images, some approaches [15, 16] try to learn part-aware objects via NeRFs. For example, Part-NeRF [15] represents an object as multiple neural radiance fields. However, the complex composition of multiple neural fields increases learning difficulty, resulting in subpar rendering outcomes and structurally insignificant information. Another state-of-the-art approach DBW [2] decomposes the scenes into multiple blocks represented by superquadric primitives [6], optimizing their parameters and UV texture maps to minimize rendering loss. Although they succeed in decomposing 3D scenes into reasonable parts, learning accurate detailed geometry and texture maps still proves challenging, or not included in the pipelines. Segmenting the scene into reasonable parts while preserving high-quality geometry details and textures is essential for good geometry reconstruction and rendering quality.
To tackle the problem, we propose integrating superquadrics and 2D Gaussians [17] to reconstruct part-aware and well-textured scenes, culminating in a hybrid representation. The hybrid representation retains the primitive property of superquadrics, where we assume each part should be a basic shape, and shapes that cannot be fitted as a superquadric should be segmented into multiple parts. On the other hand, Gaussian splatting, renowned for its superior rendering capabilities in recent studies [18, 19, 20] and efficient training, is incorporated to capture intricate texture details, speeding up the reconstruction and improving the rendering quality significantly.
@@
Specifically, in the hybrid representation, Gaussian ellipses are attached to the surface of each superquadric mesh. We adopt 2D Gaussians and share their pose parameters with the attached superquadric mesh. Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient. In this way, the parameters to optimize include global parameters controlling the shape, pose, and opacity of the hybrid block, as well as local parameters of spherical harmonic coefficients of the Gaussian ellipses. In the training stage, the parameters are optimized mainly by re-rendering loss without additional supervision. We observe that image rendering loss tends to settle into local minima during the deformation of superquadrics. To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information. This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction.
At last, to model the accurate geometry, we conduct a point-level refinement, freeing the 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes.
We summarize the contributions as follows:
@@
  • A novel hybrid representation is introduced to learn part-aware 3D scenes, where primitives model the parts and Gaussians model the textures. The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
@@
  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with novel regularizers to optimize the superquadric meshes and 2D Gaussians simultaneously.
@@
  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, working for part segmentation of objects and scenes, outperforming several state-of-the-art methods.
@@
Figure 2: Overview of the proposed method. It takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians. The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction. At last, we free the constraints of Gaussians to model the detailed geometry to get a point-level reconstruction.
II Related Works
In this section, we first introduce classic multi-view reconstruction techniques. Next, we provide an overview of the advancements in primitive-based 3D representation. Additionally, we discuss related works on surface-aligned Gaussian splatting, which inspired our hybrid representation.
II-A Multi-View Reconstruction
Before the era of deep learning, multi-view stereo-based methods [21, 22, 23, 24] adopt a standard pipeline that consisting of image feature matching, depth prediction, and fusion from different viewpoints. Recent neural approaches [25, 26, 27] combining implicit representation and neural rendering outperforms previous methods in high-fidelity reconstruction. These techniques typically represent surfaces implicitly via a multi-layer perceptron (MLP) and extract surfaces post-training using the Marching Cubes algorithm [28]. Additionally, some works [29, 30] use volumetric representations to achieve fast learning. 3D Gaussian Splatting (3D-GS) [18] was introduced as a new technique integrating point-based rendering and splitting. Due to its rendering speeds and quality never seen before, Gaussian splatting has sparked a research boom and has been applied in 3D reconstruction [17, 31, 32]. Our work can be seen as an end-to-end approach to part-aware multi-view reconstruction, whose goal is to output the 3D reconstruction of parts from calibrated multi-view images.
II-B Primitive-based 3D Representation
Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [33, 34, 35, 36, 37, 38]. The earliest classical approaches, dating back to the last century, such as Blocks World [39] and Generalized Cylinders [40] emphasize compactness as the central goal. However, these ideas struggle to handle the complexity of real-world data.
@@ II-B Primitive-based 3D Representation
A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids. These methods decompose the 3D data into various types of primitive ensembles, including cuboids [4, 41], superquadrics [42, 43, 9, 1], and convex shapes [44, 45]. Recent advancements include MonteBoxFinder [41], which utilizes a RANSAC-based [46] approach with cuboids to represent the input point cloud, combined with Monte Carlo Tree Search for optimal selection. Alternatively, EMS (Expectation, Maximization, and Switching) [1] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover these primitives from the point cloud. However, such methods often rely on perfect 3D data, rarely available in real-world scenarios. In practice, 3D data is frequently incomplete or noisy, especially when reconstructed from 2D images. This noise and missing data significantly hinder the practical application of these methods.
@@
There are only a few studies [15, 2, 16, 47] that attempted to create structure-aware 3D representations directly from images without relying on 3D supervision. For example, PartNeRF [15] uses ellipses to achieve shape abstraction, modeling each part separately using a neural radiation field. However, using ellipses makes PartNeRF more likely to decompose shapes based solely on spatial locations, which limits its ability to recover meaningful structural information. ISCO [47] and DBW [2] propose decomposing shapes using 3D superquadrics, which allows for meaningful shape decomposition. Nonetheless, the simplicity of their shape parameters struggles to capture complex geometries, leading to poor reconstruction accuracy and challenges in rendering high-quality novel view synthesis. Recently, DPA-Net [16] has made great progress on 3D shape abstraction from sparse views. However, this approach generates redundant parts and struggles with realistic texture rendering. Our approach introduces a hybrid representation to learn part-aware 3D scenes while maintaining accurate structural reconstruction and efficient, high-quality rendering capabilities.
II-C Surface-Aligned Gaussian Splatting
@@
Gaussian splatting [18] has been rapidly adopted in multiple fields because of its remarkable capability in rendering. However, the challenge of unconstrained Gaussians resulting in inaccurate geometry during reconstruction has prompted several studies [48, 49, 50, 51, 17] to align Gaussians with scene surfaces. Among them, SuGaR [48] addresses this by using flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and the actual Gaussians. GaMeS [51] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to the triangular facets of the mesh. It allows for mesh-like adaptation and modification of Gaussian components. Similarly, Gao et al. [50] propose a mesh-based Gaussian Splatting to achieve large-scale deformation effects on objects. NeuSG [49] combines 3DGS with implicit SDF networks, optimizing the Gaussian components and the network parameters jointly. Recently, 2DGS [17] proposes using 2D Gaussians instead of 3D Gaussians, leveraging 2D surface modeling to significantly enhance the geometric quality of scenes. Inspired by them, our work proposes a representation that combines 2D Gaussians with superquadrics, ensuring that the Gaussians conform to the surface of the superquadrics.
III Method
Given a set of calibrated multi-view images, we aim to learn the part-aware 3D geometry and texture information based on Gaussian splatting. We observe that a part-aware scene can intuitively be regarded as a composition of multiple basic shapes (geometric primitives). Therefore, we design a primitive-based hybrid representation to describe the 3D scene, as introduced in Section III-A. In Section III-B, this hybrid representation can be optimized to fit primitives as different parts, as stage 1. In stage 2, we refine the positions of 2D Gaussians to model the precise geometry of the original shape.
III-A Hybrid Representation
As shown in Fig. 2, the proposed hybrid representation combines superquadrics with 2D Gaussian ellipses in a compact form. Superquadrics are efficient for modeling 3D scenes due to their primitive properties, but they fall short in capturing detailed geometry and texture information. Conversely, the 3D/2D Gaussian representation [18, 17] excels in photorealistic rendering but is cumbersome and lacks structural awareness.
@@ III-A Hybrid Representation
To leverage the strengths of both representations, we attach 2D Gaussian ellipses to the surface of the superquadric mesh. This approach retains the superquadric’s advantage of effectively parsing a 3D scene into distinct parts. At the same time, the Spherical Harmonics of 2D Gaussians enable the rendering of complex textures by Gaussian splatting, addressing the limitations in texture learning. By sharing the pose parameters between superquadrics and 2D Gaussians, we achieve a more efficient representation.
III-A1 The Parametric Model
@@
The proposed hybrid model is parametric, controlled by primitive parameters and Gaussian parameters. Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e. blocks, in the first stage. Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$. Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. The hybrid blocks are initialized using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized through a multi-view rendering loss to learn the scene.
@@
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its shape is defined by the superquadric mesh. As introduced in [6], a superquadric comprises a series of parametric surfaces controlled by several continuous parameters. More concretely, there are 2 parameters $\epsilon_{1},\epsilon_{2}$ defining the superquadric mesh and 3 anisotropy parameters $s_{1},s_{2},s_{3}$ scaling the 3D axes. Similar to the position of each vertex of an icosphere, we can acquire the 3D position of the superquadric vertex by:
$\scriptstyle\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
@@
where $\theta$ and $\varphi$ are the azimuth and elevation defined in the spherical coordinate system. The shape and scale parameters are optimized during the training stage, thus controlling the block deformation to learn part-aware information.
@@
Pose Parameters. The initial hybrid blocks are randomly sampled and disorganized in the space, so it is necessary to align their poses to fit within the 3D scene. The vertex position acquired in Eq. 1 is in the local coordinate system of the hybrid block, so we need to use the pose parameters to transform the position into the world space. We define the pose parameter of the $i$-th hybrid block as the combination of rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. Hence, the vertex position in the world space is calculated by: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $i,j$ denotes the indexes of block and vertex respectively. In general, the Gaussians placed on the superquadric surface also have their individual poses, which causes a large amount of parameters. To address this issue, we propose to share the pose of the superquadric with the Gaussians. Given a posed superquadric, we directly determine the center points of the Gaussians through uniform sampling on the triangular faces and let the normal of 2D Gaussians be the same as the faces. Inspired by GaMeS [51], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on the vertices’ positions. For a given triangular face $V$ and its three vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define the orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$. The first vector $r_{1}$ is aligned with the normal vector of $V$ to ensure the Gaussians align with the triangular surface. The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
@@
where ${ort}$ represents the orthonormalization step in the Gram-Schmidt [52]. For the scaling parameters of the 2D Gaussians, we use:
@@
where $s_{v}^{2}=c||m-v_{1}||$, and $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of $n$ Gaussians on each triangular face.
This design avoids the additional cost of optimizing the pose and covariance matrix of Gaussians, making it more efficient for representing the scene. The pose parameters are optimized via gradient descent so that the hybrid blocks can fit the scene.
@@
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, in practice, a typical scene does not contain exactly $M$ blocks, so we retain only the meaningful ones. Therefore, we utilize a learnable parameter $\tau_{i}$ to represent each block’s opacity. In the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
@@
Texture Parameters. The texture information is modeled using 2D Gaussians positioned on the surface of the superquadric mesh. Spherical harmonics control the texture of each Gaussian ellipse. Therefore, we define the coefficients of the spherical harmonics as texture parameters and optimize them for rendering view-dependent images.
III-A2 Rendering
With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [18, 17]. Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image. Initially, the screen space determines the bounding box for each Gaussian. Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane. Finally, volumetric alpha compositing [53] is utilized to integrate the alpha-weighted RGB values for each pixel.
To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (5)
where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse. Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians. Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (6)
Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located. The calculated alphas are sorted according to their depths from the image plane. Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (7)
This process is differentiable and can optimize the hybrid representation through gradient descent, and the optimization/learning is introduced in Sec. III-B.
III-B Optimizing the Hybrid Representation

### v2

# PartGS: Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

@@ Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes. However, human perception typically understands 3D objects at a higher level as a composition of parts or structures rather than points or voxels. Representing 3D objects or scenes as semantic parts can benefit further understanding and applications. In this paper, we introduce PartGS, part-aware 3D reconstruction by a hybrid representation of 2D Gaussians and Superquadrics, which parses objects or scenes into semantic parts, digging 3D structural clues from multi-view image inputs. Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time. Our method simultaneously optimizes superquadric meshes and Gaussians by coupling their parameters within our hybrid representation. On one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes. On the other hand, 2D Gaussians capture complex texture and geometry details, ensuring high-quality appearance and geometry reconstruction. Our method is fully unsupervised and outperforms existing state-of-the-art approaches in extensive experiments on DTU, ShapeNet, and real-life datasets.
1 Introduction
@@
3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which differ from human perception. Humans understand 3D scenes as different semantic parts [26]. For instance, when observing a scene, we mentally construct high-level structural information, such as scene graphs, rather than point clouds or detailed geometry. Thus, we propose to learn part-aware scene reconstruction to decompose scenes into different individual semantic shapes or parts, facilitating tasks including physical simulation, manipulation/editing, 3D content generation, etc.
@@
There are prior methods [38, 51, 2, 28, 20, 44, 22, 23, 36, 42, 48] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios. While NeRF [25] has demonstrated great potential in reconstructing textured 3D scenes from multi-view images, some approaches [37, 49] try to learn part-aware objects via NeRFs. In particular, Part-NeRF [37] represents an object as multiple neural radiance fields. However, the complex composition of multiple implicit fields increases learning difficulty, resulting in subpar rendering quality and insufficient semantic decomposition. Another state-of-the-art approach DBW [27] decomposes scenes into blocks represented by superquadric primitives [2], optimizing their parameters and UV texture maps by minimizing rendering loss. Although DBW effectively decomposes 3D scenes into reasonable parts, learning accurate part-aware geometry and appearance still proves challenging as shown in Fig. 1.
To tackle the problem, we propose PartGS, a part-aware hybrid representation that integrates 2D Gaussians [16] and superquadrics [2] to reconstruct scenes that are both well-textured and geometrically accurate. There are previous approaches [15, 13, 39] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it. Their purpose is to get both representations, while our goal is to get part segmentation and reconstruction at the same time. Our method directly learns superquadric mesh and Gaussians simultaneously by coupling parameters between them while constraining each part to be represented as a single superquadric. The rendering capability of coupled Gaussians drives the optimization of hybrid representation, enabling the use of their convexity to assist in the decomposition of shapes. The unsupervised training of part-aware reconstruction is built upon the assumption that each functional part of most objects should be a basic shape and can be represented by a superquadric. On the other hand, Gaussian splatting, renowned for its superior rendering capabilities and efficient training[18, 11, 24], is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality significantly. In the hybrid representation, 2D Gaussians are attached to the surface of superquadrics, with their pose and shape parameters coupled to the corresponding superquadric mesh. Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient. The optimized parameters include global ones controlling the overall shape, pose, and opacity of hybrid blocks, and local parameters for spherical harmonic coefficients of the Gaussians.
@@
During training, the parameters are optimized by rendering loss without additional supervision. We observe that image rendering loss tends to settle into local minima during the optimization of superquadrics. To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information. This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction. At last, to model the accurate geometry, we conduct a point-level refinement, freeing 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes. Contributions are summarized as follows:
@@
  • A novel hybrid representation is introduced to learn part-aware textured 3D scenes. The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
@@
  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with superquadric meshes and 2D Gaussians optimized simultaneously by novel regularizers.
@@
  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, enabling applications of part augmentation and physical interactions.
@@
Figure 2: Overview of PartGS. PartGS takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians. The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction. Then, we free the constraints of Gaussians to model detailed geometry to get a point-level reconstruction. The last column demonstrates one application of PartGS: physics simulations, where one part is picked up from the scene and thrown forward. Please refer to the supplementary materials for more video demos.
2 Related Works
2.1 Primitive-based 3D Representation
@@
Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [43, 50, 14, 9, 41, 35]. The earliest classical approaches, dating back to the last century, such as Blocks World [32] and Generalized Cylinders [3] emphasize compactness as the central goal. A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids. These methods decompose the 3D data into various types of primitive ensembles, including cuboids [38, 31], superquadrics [29, 30, 44, 22], and convex shapes [8, 10]. MonteBoxFinder [31], uses a RANSAC-based [12] with cuboids to represent point clouds, combined with Monte Carlo Tree Search for optimal selection. Alternatively, EMS [22] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover primitives. These methods rely on perfect 3D data, which is rarely available due to incompleteness or noise.
@@
Few studies [37, 27, 49, 1] have attempted to create structure-aware 3D representations directly from images. PartNeRF [37] uses ellipses with neural radiance fields, but this spatial-based decomposition limits its ability to capture meaningful structural information. ISCO [1] and DBW [27] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation. However, their simple shape parameters struggle to capture complex geometries, leading to poor geometry and appearance reconstruction. DPA-Net [49] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering. Our approach introduces a hybrid representation that learns part-aware 3D scenes while ensuring accurate structural reconstruction and efficient, high-quality rendering.
2.2 Mesh-based Gaussian Splatting
@@
Gaussian splatting(GS) [18] has been rapidly adopted in multiple fields due to its remarkable rendering capability. Several studies [15, 7, 13, 39] attempt to align Gaussians with mesh surfaces for easier editing and animation. Among them, SuGaR [15] uses flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and actual Gaussians. GaMeS [39] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to triangular facets of the mesh. Similarly, Gao et al. [13] proposes a mesh-based GS to achieve large-scale deformation effects on objects. Recently, 2DGS [16] proposes 2D Gaussians for surface modeling and significantly enhances the geometric quality. Inspired by them, we propose a representation that combines 2D Gaussians with superquadrics. A key distinction between our method and previous mesh-based GS approaches is that these methods require first reconstructing the scene’s mesh and then binding Gaussians to the mesh surface, which results in a non-continuous mesh. In contrast, our method directly optimizes the mesh through a rendering loss, enabling part-aware mesh reconstruction.
3 Method
Given a set of calibrated multi-view images, our goal is to learn part-aware 3D geometry and appearance. A part-aware reconstruction is often represented as a composition of multiple basic shapes (geometric primitives). We propose a primitive-based hybrid representation in Sec. 3.1. In Sec. 3.2, we demonstrate how this hybrid representation is optimized to fit primitives as parts (block-level). In order to further detail the geometry of each part, the parameters of Gaussians are further refined to get a point-level reconstruction.
3.1 Hybrid Representation
@@
As shown in Fig. 2, the proposed hybrid representation combines superquadrics with 2D Gaussians in a compact form. To leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes. This representation retains the superquadric’s ability to effectively parse a 3D scene into distinct parts. Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of previous work [22, 31, 27, 48]. Sharing pose parameters between superquadrics and 2D Gaussians further enhances the representation’s efficiency.
@@
The proposed hybrid model is parametric, controlled by both primitive and Gaussian parameters, which are optimized simultaneously. Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e. blocks. Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$. Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. The hybrid blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized via differentiable rendering to parse the target scene.
@@
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2]. More concretely, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define the superquadric mesh, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes. Similar to the positioning of vertexes in an icosphere, the 3D position of a superquadric vertex can be obtained by:
$\footnotesize\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
@@
where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate, respectively. The shape and scale parameters are optimized during training, thus controlling the block deformation to learn part-aware geometry.
Pose Parameters. The pose of the $i$-th hybrid block is defined by its rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. The vertex position from Eq. 1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices. In prior methods such as SuGaR [15] and GaMeS [39], Gaussians are placed on the reconstructed mesh, and they have individual poses and shapes for better appearance modeling. Differently, in our method, the role of Gaussians is to construct a differentiable rendering that bridges images and superquadrics. To achieve this, we couple the parameters of Gaussians with superquadrics.
@@
Given a posed superquadric, centers of Gaussians are determined by uniformly sampling on triangular faces, and poses of Gaussians are determined by face vertices. Inspired by GaMeS [39], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on vertices’ positions. For a given triangular face $V$ and its vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$. The vector $r_{1}$ is aligned with the normal vector of $V$ to ensure Gaussians align with the triangular surface. The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
@@
where ${ort}$ represents the orthonormalization in Gram-Schmidt [4]. For the scaling of 2D Gaussians, we use:
@@
where $s_{v}^{2}=c||m-v_{1}||$, $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of Gaussians on each triangular face. These designs eliminate the need to learn the geometry of Gaussians, focusing solely on their texture, thereby enhancing the efficiency of the representation.
@@
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, in practice, a typical scene does not contain exactly $M$ blocks. Therefore, only the meaningful blocks are retained. To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity. During the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
@@
Texture Parameters. The texture is modeled using 2D Gaussians positioned on the surface of the superquadrics. Spherical harmonics of each Gaussian control the texture and are optimized for rendering view-dependent images [18].
3.2 Optimizing the Hybrid Representation